BaconLover:

Out of the 120 students at a summer camp, 72 signed up for canoeing. There were 23 students who signed up for trekking, and 13 of those students also signed up for canoeing. Use a two-way table to organize the information and answer the following question: Approximately what percentage of students signed up for neither canoeing nor trekking? (1 point) 10% 12% 38% 32%

BaconLover:

@undeadknight26

Yinshy:

well first what do you think and why?

Yinshy:

@BaconLover

BaconLover:

well i was gonna say 12

Yinshy:

may i ask why you think that?

BaconLover:

only because i did 72+23+13 and got 108 btu when i did 120-108 i got 12

Yinshy:

oooo close but that 13 is apart of the 23 because it says 13 of the 23 also signed up for canoeing

Yinshy:

so what do you think now?

BaconLover:

so what your saying is that i was close to the answer? so im gonna say 10 perccent then

Yinshy:

wait my logic is flawed sorry

Yinshy:

one sec

Yinshy:

lemi get the math wiz in here

Yinshy:

@Ultrilliam

Yinshy:

like i said im a writer not a math wiz xD

Yinshy:

but now as i think of it 12% would be correct

BaconLover:

so my math was right all along ? -_-

Yinshy:

yes ^^;

BaconLover:

good

Yinshy:

if you any more questions tag me

BaconLover:

ok will do

undeadknight26:

How did you guys get 12%? Wasn't it 72 + 23 = 95 - 13 = 82. 120 - 82 = 38/120 = .32?

undeadknight26:

Or did I go wrong?

BaconLover:

your math seems more real thank you and i got 12 by doing 72+23+13=108 so i did 120-108 and got 12

undeadknight26:

No you have to subtract 13 from the sum of 72 and 23. If you add it you'll be counting the same person twice.

BaconLover:

oh

Warriorz13:

Babe

BaconLover:

yes?

Warriorz13:

Hi

BaconLover:

lol are you gonna spam my math question?

Warriorz13:

Maybe

BaconLover:

badddddd